OpenStudy (anonymous):

The equation of line AB is (y-3) = 5 (x - 4). What is the slope of a line perpendicular to line AB? A) -5 B) 5 C) -1/5 D) 1/5 My answer was C, is this right?

OpenStudy (jdoe0001):

what's the slope of the line AB anyway?

OpenStudy (anonymous):

You're right. The slope of AB is 5 and the perpendicular is - 1/5

OpenStudy (anonymous):

Thanks! Can you help me with another one? The equation of line CD is y = -2x - 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5). A) y = -2x + 13 B) y = -1/2 x + 7 C) y = -1/2 x + 3 D) y = − 2x − 3

OpenStudy (anonymous):

\[y = mx +b\] Take the slope intercept formula and plug in what we know We know that the slope is -2 (m), x is 4, and y is 5. So plug those number into the equation and solve for b \[5 = -2(4) + b\] Simplify \[5 = -8 + b\] add 8 to both sides \[13 = b\] and because we are looking for a parallel line, the slopes are the same. So, y = -2x + 13

OpenStudy (anonymous):

omg you are a life saver :) just one more pleeeassee :) The equation of line WX is y = -2x - 5. Write an equation of a line perpendicular to line WX in slope-intercept form that contains point (-1, -2). A) y = (1/2)x + 3/2 B) y = (-1/2)x + 3/2 C) y = (1/2)x - 3/2 D) y = (-1/2)x - 3/2

OpenStudy (anonymous):

I got C